In 15 experiments studying electrical behavior of single cells 11 use microelectrodes

made of metal and the other 4 use microelectrodes made from glass tubing. If 2 of the

experiments are to be terminated for financial reasons and they are selected at random.

What are the probabilities that

a. Neither uses microelectrodes made from glass tubing?

b. Only one use microelectrodes made from glass tubing?


1
Expert's answer
2022-04-07T07:16:08-0400

"a:\\P=\frac{C_{11}^{2}}{C_{15}^{2}}=\frac{55}{105}=\frac{11}{21}\\b:\\P=\frac{C_{11}^{1}C_{4}^{1}}{C_{15}^{2}}=\frac{44}{105}"